OpenStudy (anonymous):

Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form. A. (y + 1) = 3(x – 2) B. (x + 2) = 3(y – 1) C. (x – 2) = 3(y + 1) D. (y – 1) = 3(x + 2)

OpenStudy (anonymous):

D !

OpenStudy (anonymous):

in order to determine the equation, we use this formula y-y1 = m(x-x1) where m = slope of the line and (x1,y1) = points throught which the line passes

OpenStudy (anonymous):

only d verifies for the point (-2, 1)

OpenStudy (anonymous):

0=0

OpenStudy (anonymous):

d is correct. I got it wrong thanks to @Dmb1215